Sol Summary – May 21st

Crew 281 Sol Summary Report 21 May 2023
Sol: 7
Summary Title: Radio contacts, robotic arm and drone drills
Author’s name: Ritupriya Patil (XO) and Megan Kane (Commander)
Mission Status: Nominal
Sol Activity Summary:
Today we did our EVA around the hab with a three-tier objective. First was to test Ana’s (Crew Scientist) Pegasus scoop, a robotic rock sampling device, it passed the testing and has now demonstrated TRL 6. The second objective was to perform a drone test to simulate emergency assistance and deliver a lightweight medical kit to an injured astronaut. This was also successful and Ritupriya (XO, EVA Lead) was able to fly the drone and deliver a 30 grams payload using her DJI Mavic Mini to the astronaut in need during the simulation. This was a 60 m long distance at 10 m altitude covered by the drone and the payload was received in 1 minute and 19 seconds. There will be more drills to iron out some of the kinks in the procedures. The third objective was to do a quick check on the ham radio antenna and ensure there was no damage due to the usual “Martian storms” that hit MDRS in the late afternoons. After inspection there were no issues reported and the antenna survived the high winds.
For lunch KC made fusilli pasta with freeze-dried veggies, black bean burger “meatballs” and alfredo pasta. After that filling lunch, Megan (Commander) tempted us with her freshly baked and piping hot cinnamon rolls which were just scrumptious! We did notice the difference in baking times possibly due to higher altitude!
Post lunchtime lull died soon since we immersed ourselves in mid-mission summary reports and I started my daily reports for our mission. Meanwhile Rachel (HSO) was able to establish contact on the ham radio setup with NR1Z.
In the green hab, Megan (Commander) did some pruning and watered the plants. KC (Crew Engineer) did routine checks on hab operations – batteries, tunnels, toilets and water levels. Ana cleaned up her robotic scoop in the RAM after a successful mission today.
In the science dome it was quiet today, cacao seeds still doing well.
For dinner we had corn tortillas with freeze-dried – guacamole, bean burger filling, beef and some canned tuna. It rained in the late afternoon and there were storms, but the antenna stayed upright.
Look Ahead Plan: Tomorrow we will do an EVA for 3 and cover Pooh’s Corner for second scanline measurement at the site and take detailed vertical and diagonal images and videos of the scanline to confirm the weathering grade and discontinuities of the rocks. On our way back we will stop at Marble Ritual for images of scanlines that were measured on earlier EVAs.
Anomalies in work: None
Weather: Nominal
Crew Physical Status: Nominal
EVA: Completed 1 today

Reports to be filed:
Greenhab Report
Ops Report
Journalist Report
Photo of the Day

Support Requested:
Mission support is requested in case of storms if the antenna needs to come down or be retrieved.

Copyright © The Mars Society. All rights reserved. | Main Site